In a recent year, the average annual earnings for a worker with a college degree were $37,000. This represents an 85% increase over the average annual earnings for a high school graduate that year. Find the average annual earnings for a high school graduate that year.
Answer by Theo(13342)


let x = the average annual salary of the high school graduate that year.

the average annual salary of the college graduate was 85% more than that in the same year.

therefore, x + .85 * x = 37,000

factor out the x to get (1 + .85) * x = 37,000

combine like terms to get 1.85 * x = 37,000

divide both sides of this equation by 1.85 to get x = 37000 / 1.85 = 20,000.

the high school graduate earned 20,000.

85% of this is equal to 20,000 * .85 = 17,000

the college graduate earned 17,000 more than 20,000 which is equal to 37,000.

your solution is that the average annual salary of a high school graduate was 20,000 that year.